OpenStudy (anonymous):

Quadratic functions, I need to solve this problem: -x^2 + 2x - 2 = 0 I think in this case, I cannot factor or use the quadratic formula. I think it's supposed to be solved using the square root method and I'm having difficulty getting an answer.

OpenStudy (anonymous):

why can you not use the quadratic formula?

OpenStudy (anonymous):

I tried, it doesn't add up properly.

OpenStudy (anonymous):

@robz8

OpenStudy (anonymous):

the quadratic formula will always work if you do it properly. it will not always be pretty, but it will give you the correct roots each time

OpenStudy (lucaz):

true

OpenStudy (anonymous):

okay; x= - b ±√ b^2 – 4ac / 2a a = 1 b = -2 and c =2 give me a sec to finish it up!

OpenStudy (anonymous):

in this equation though, you will get imaginary roots, since the equation does not cross the x axis

OpenStudy (jhannybean):

\[\large y=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]b= 2, a= -1 c=-2\[\large y= \frac{-2 \pm \sqrt{(2)^2-4(-1)(-2)}}{2(-1)}\]\[\large y= \frac{-2 \pm \sqrt{4-8}}{2(-1)}\] It will not work.

OpenStudy (anonymous):

I'm also having difficulty, it's supposed to have an imaginary number as the answer, but that's not coming up.

OpenStudy (anonymous):

the imaginary roots are x = 1 - i and x = 1+i

OpenStudy (jhannybean):

Yes.

OpenStudy (anonymous):

how did you solve that though?

OpenStudy (anonymous):

@robz8 @Jhannybean

OpenStudy (jhannybean):

Just a minute.

OpenStudy (anonymous):

take your time, im solving it as well!

OpenStudy (jhannybean):

\[\large -(-x^2 + 2x -2) = 0 \]multiply the negative in \[\large x^2-2x+2 =0 \] complete the square \[\large x^2-2x =- 2\] take 2 divide that by half and square it. Then take that answer, it will become your new "c" Also, add it to the 2 on the other side. \[\large x^2-2x+1 = -2+1\]simplify \[\large x^2-2x+1 =-1\] you can write the squared form on the left hand side. \[\large (x-1)^2=-1\]square root both sides to solve for x. \[\large \sqrt{(x-1)^2}= \sqrt{-1}\]the square root of a squared results in an absolute value\[\large |x-1| = i\] now you will have two solutions. \[\large x-1 = i\] and \[\large x-1=-i\] isolate x for both these solutions. \[\color{orange}{x=i+1 , x=1-i}\]

OpenStudy (anonymous):

Wow, my lesson does not go into this kind of detail. Thanks, this makes so much more sense than what I was doing!! @Jhannybean

OpenStudy (jhannybean):

Are you able to follow step by step?

OpenStudy (anonymous):

yes!

OpenStudy (jhannybean):

great :D

OpenStudy (anonymous):

thanks (: @Jhannybean

OpenStudy (jhannybean):

no problemo :P

OpenStudy (anonymous):

@Darrius